Angle 1 and angle 2 form a linear pair.Find the measure of each if the measure of the obtuse is four times the measure of the acute angle.
Answer by stanbon(75887)

Angle 1 and angle 2 form a linear pair.Find the measure of each if the measure of the obtuse is four times the measure of the acute angle.
------
Equations:
x + y = 180 degrees
x = 4y
-----
Substitute and solve for "y":
4y+y = 180
5y = 180
y = 36 degrees
x = 4y = 4*36 = 144 degrees
============================
Cheers,
Stan H.
=============